OpenStudy (anonymous):

The coach of a cricket team buys 3 bats and 6 balls for Rs .3900.Later,she buys another bat and 3 more balls of the same kind for Rs 1300.Represent this situation algebraically.

OpenStudy (diyadiya):

Let Number of bats be x & Number of Balls be y 3 bats and 6 balls for Rs .3900 3x + 6y =3900 What will be the next equation?

OpenStudy (anonymous):

1st equation wch u types i knw bt i can't frame d 2nd one.

OpenStudy (diyadiya):

Later,she buys another bat and 3 more balls tell me how many bats & balls she has now?

OpenStudy (anonymous):

i dnt knw

OpenStudy (anonymous):

is it x+2y=1300

OpenStudy (diyadiya):

Where did you get that 2 from/?

OpenStudy (anonymous):

dts wht i wrote in ma notebook

OpenStudy (anonymous):

x+3y=1300 is it correct?

OpenStudy (diyadiya):

Later,she buys another bat and \(\Large 3 \) more balls of the same kind for Rs 1300

OpenStudy (anonymous):

is it x+3y=1300

OpenStudy (anonymous):

is it diya?

OpenStudy (diyadiya):

Yes! :)

OpenStudy (anonymous):

k

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

thx :)

OpenStudy (diyadiya):

No problem :)